I have to use either elimination or substitution to find the solutions and show my work. Then I will have to graph it. The graphin is not my problem on this one. Trying to find the intercepts and slopes is. I have to have 3 points. The equation is :
5x + 6y = 1 and x + 7y = 2
This is what I did but I don't think its right and If it is, i don't know where to go from there to get the other points.
x=2-7y
5(2-7y) + 6y=1
10-35y = 6y=1
-29y= -9
y= 9/29 or 0.310344828
x=2-7(9/29)
x= -5/29 or -0.172413793

Answer by stanbon(75887)

This is what I did but I don't think its right and If it is, i don't know where to go from there to get the other points.
x=2-7y
5(2-7y) + 6y=1
10-35y + 6y=1
-29y= -9
y= 9/29 or 0.310344828
x=2-7(9/29)
x= -5/29 or -0.172413793
-------------
YOU ARE CORRECT
I graphed it with a graphing calculator.
Your x,y values are the point of intersection.
==================
Cheers,
Stan H.